### v1

# Empowering Snapshot Compressive Imaging: Spatial-Spectral State Space Model with Across-Scanning and Local Enhancement

Abstract
Snapshot Compressive Imaging (SCI) relies on decoding algorithms such as CNN or Transformer to reconstruct the hyperspectral image (HSI) from its compressed measurement. Although existing CNN and Transformer-based methods have proven effective, CNNs are limited by their inadequate modeling of long-range dependencies, while Transformer ones face high computational costs due to quadratic complexity. Recent Mamba models have demonstrated superior performance over CNN and Transformer-based architectures in some visual tasks, but these models have not fully utilized the local similarities in both spatial and spectral dimensions. Moreover, the long-sequence modeling capability of SSM may offer an advantage in processing the numerous spectral bands for HSI reconstruction, which has not yet been explored. In this paper, we introduce a State Space Model with Across-Scanning and Local Enhancement, named ASLE-SSM, that employs a Spatial-Spectral SSM for global-local balanced context encoding and cross-channel interaction promoting. Specifically, we introduce local scanning in the spatial dimension to balance the global and local receptive fields, and then propose our across-scanning method based on spatial-spectral local cubes to leverage local similarities between adjacent spectral bands and pixels to guide the reconstruction process. These two scanning mechanisms extract the HSI’s local features while balancing the global perspective without any additional costs. Experimental results illustrate ASLE-SSM’s superiority over existing state-of-the-art methods, with an inference speed 2.4$\times$ faster than Transformer-based MST and saving 0.12 (M) of parameters, achieving the lowest computational cost and parameter count.
1 Introduction
Figure 1: PSNR-FLOPs comparisons of ASLE-SSM and previous End-to-End Transformer-based SOTA methods.
Hyperspectral images (HSIs) contain a large number of spectral bands to store rich detail information and thus are widely used in various fields such as object tracking ot_1 ; ot_2 ; ot_3 , remote sensing chen2017denoising ; he2020non ; yuan2018hyperspectral , super-resolution dong2021model ; sr_2 ; sr_3 , and medical image processing mi_1 ; mi_2 ; mi_3 . Traditional imaging systems used for collecting HSIs involve scanning along the spatial or spectral dimensions but this imaging process is time-consuming and can not effectively capture dynamic objects. Recently, snapshot compressive imaging (SCI) systems have been developed sci_1 ; sci_2 as an alternative method for obtaining HSIs at video rate. SCI systems capture multiple image bands in a single snapshot, compressing the information of HSIs into a 2D measurement. sci_3 ; sci_4 ; sci_5 ; sci_6 . Among these systems, Coded Aperture Snapshot Spectral Imaging (CASSI) has emerged and become mainstream gap_tv ; icme_sci_1 . CASSI utilizes a physical mask to modulate signals from HSIs, shifting these modulated images to different spatial locations on the detector plane via a disperser mst ; icme_sci_2 . However, the performance of the CASSI system heavily depends on the efficiency of the reconstruction algorithm.
In recent years, numerous reconstruction algorithms have been proposed, including model-based methods and deep learning-based methods. Traditional model-based methods typically attempt to model the inverse process of image reconstruction, using manually designed priors to derive the solution to an optimization problem. Although these methods are highly interpretable, they often require manual parameter tuning and tend to be slow. With the success of deep learning methods in various visual tasks, CNN and Transformer-based methods have been developed and applied to HSI reconstruction, achieving higher image reconstruction quality and speed than traditional algorithms. The CNN-based methods lambda ; tsa_net have an advantage in extracting local features from images, but they lack the ability to model long-range dependencies effectively. Transformer-based methods mst ; mst_pp ; dauhst ; cst realize a global receptive field, resulting in superior performance compared to CNNs but higher computational costs. Although some mechanisms can mitigate this issue, they do not fundamentally change the quadratic complexity problem caused by the attention mechanism in the Transformer architecture.
Figure 2: The flowchart of the proposed ASLE-SSM. The framework is arranged by adopting a U-shaped structure. (a) Spatial-Spectral SSM Block. This module mainly consists of a spatial SSM and a spectral SSM, depicted in the sub-flowchart (b) and (c), respectively. (d) The specific method of LE-SSM in the Spatial SSM Module. Scanning directions are divided into global and local branches. (e) The illustration of our Across-Scanning Mechanism. The scanning process is demonstrated by labeling the order of each data point.
The limitation of Transformers in spectral SCI remains unresolved. However, in the realm of large language model and RGB image vision, an advanced solution for this problem exists, known as Mamba gu2023mamba . Through data-dependent selective SSMs with parallel scanning and hardware-wise strategies, Mamba establishes a global receptive field and achieves near-linear complexity. Mamba-based methods have been shown to outperform previous methods in various tasks due to their excellent long sequence modeling capabilities and lower computational costs mamba_survey ; vmamba2 ; mambair ; lightm-unet .
Subsequently, many works have begun to explore the application and performance of the Mamba model in various visual tasks. Representative works, such as the visual state space model shi2024vmambair and vision Mamba vision_mamba , have demonstrated Mamba’s superiority for image processing, particularly in obtaining fewer parameters and faster speed. However, most Mamba-based models only conduct the scanning process along the spatial dimension, lacking cross-channel interaction, despite the importance of spectral information for HSI reconstruction. Recently, some Mamba-based methods have made progress in developing an SSM mechanism along the channel (spectral) dimension for certain visual tasks s2mamba ; huang2024spectral , achieving better results compared to scanning only along the spatial dimension. However, this scanning method does not consider the inherently local features of HSIs in both spatial and spectral dimensions.
In this paper, we introduce ASLE-SSM, a spatial-spectral SSM with across-scanning and local enhancement, to improve HSI reconstruction in SCI. Our ASLE-SSM aims to strike a balance between global and local information utilization in HSIs. ASLE-SSM comprises two main components: spatial SSM and spectral-spatial cross SSM. The spatial SSM employs a combination of global and local scanning methods, enhancing row-by-row scanning with both global and local modeling. Meanwhile, the spectral-spatial cross SSM incorporates an innovative across-scanning mechanism, leveraging spatial-spectral local cubes to enhance the module’s ability to capture local details. Overall, our contributions are summarized as follows:
  • We present a novel spectral-spatial state space model for HSI reconstruction, which can simultaneously integrate information from both spectral and spatial dimensions to adequately model spatial context and spectral features.
  • We propose an across-scanning method that integrates local features from both spatial and spectral dimensions within a single scan. Additionally, we introduce a local scanning approach within the spatial SSM to further enhance local feature extraction.
  • Our network outperforms state-of-the-art (SOTA) methods on simulation and real datasets, delivering visually superior results. Importantly, it boasts the minimum parameter count, lowest computational cost, and fastest inference speed among current SOTA methods.
Figure 3: A Schematic diagram of the coded aperture snapshot spectral imaging (CASSI) system. The 3D HSI cube is first modulated by a physical mask and then the modulated data cube is sheared by a disperser. The 2D-coded measurement thus includes the information on the HSI data cube.
2 Preliminaries and Related Works
2.1 CASSI System
CASSI consists of a hardware encoder and a software decoder. The encoder compresses a 3D data cube into a 2D snapshot measurement and the decoder reconstructs HSIs from the measurement. An imaging process CASSI system is illustrated in Fig. 3. The 3D HSI data cube is denoted as $\mathbf{F}\in\mathbb{R}^{H\times W\times N_{\lambda}}$ where H, W, and $N_{\lambda}$ represent the height, width, and number of wavelengths of the 3D HSI, respectively. A physical mask $\mathbf{M}\in\mathbb{R}^{H\times W}$ is used to modulate the 3D HSI $\mathbf{F}$ as:
$\mathbf{F^{{}^{\prime}}}(:,:,n_{\lambda})=\mathbf{F^{{}^{\prime}}}(:,:,n_{\lambda})\odot\mathbf{M},$ (1)
where $\odot$ represents the element-wise product and $\mathbf{F^{{}^{\prime}}}\in\mathbb{R}^{H\times W\times N_{\lambda}}$ is the modulated 3D HSIs. After the modulated cube passes the disperser, $\mathbf{F^{{}^{\prime}}}$ is tilted and considered to be sheared along the y-axis. We use $\mathbf{F^{{}^{\prime\prime}}}\in\mathbb{R}^{H\times(W+d(N_{\lambda}-1))\times N_{\lambda}}$ to denote the tilted cube and assume $n_{\lambda}$ to be the reference wavelength. That is, image $\mathbf{F^{{}^{\prime}}}(:,:,n_{\lambda_{c}})$ is not sheared along the y-axis. Then we have:
$\mathbf{F^{{}^{\prime\prime}}}(u,v,n_{\lambda})=\mathbf{F^{{}^{\prime}}}(x,y+d(\lambda_{n}-\lambda_{c}),n_{\lambda}),$ (2)
where $(u,v)$ indicates the coordinate system on the detector plane, and $\lambda_{n}$ is the wavelength of channel $n_{\lambda}$. Here, $d(\lambda_{n}-\lambda_{c})$ signifies the spatial shifting for channel $n_{\lambda}$. Finally, the 2D compressed measurement $\mathbf{Y}\in\mathbb{R}^{H\times[W+d(N_{\lambda}-1)]}$ can be obtained by
$\mathbf{Y}=\sum_{n_{\lambda}=1}^{N_{\lambda}}\mathbf{F^{{}^{\prime\prime}}}(:,:,n_{\lambda})+\mathbf{N},$ (3)
where $\mathbf{N}\in\mathbb{R}^{H\times[W+d(N_{\lambda}-1)]}$ represents the noise of CASSI caused by the sensing detector.
2.2 State Space Models
Recently, researchers have made significant progress in state space models (SSMs), leading to the emergence of various methods and architectures based on SSMs. Among them, the structured state space model (S4) s4gu2021 becomes a pioneer work for its advantage in modeling long-range dependencies. S4 maps a 1D sequence $x(t)\in\mathbb{R}\rightarrow y(t)\in\mathbb{R}$ by an implicit latent state $h(t)\in\mathbb{R}^{N}$ and decodes it into an output sequence, which can be formulated as:
$h^{\prime}(t)=\mathbf{A}h(t)+\mathbf{B}x(t),$ (4)
$y(t)=\mathbf{C}h(t)+\mathbf{D}x(t),$
where $N$ is the state size, ${\rm\textbf{A}}\in\mathbb{R}^{N\times N}$, ${\rm\textbf{B}}\in\mathbb{R}^{N\times 1}$, ${\rm\textbf{C}}\in\mathbb{R}^{1\times N}$, and ${\rm\textbf{D}}\in\mathbb{R}$. After that, a discretization process is adopted to integrate Eq. (4) into the deep learning algorithm, using the zero-order hold (ZOH) rule, which is defined as follows:
$\overline{\rm\textbf{A}}$ $={\rm exp}({\rm{\Delta\textbf{A}}}),$ (5)
$\overline{\rm\textbf{B}}$ $=({\rm{\Delta\textbf{A}}})^{-1}({\rm exp(\textbf{A})}-\textbf{I})\cdot{\rm\Delta\textbf{B}},$
where $\overline{\rm\textbf{A}}$ and $\overline{\mathbf{B}}$ represent the discretized forms of the parameters $\mathbf{A}$ and $\mathbf{B}$, respectively. $\rm\Delta$ is denoted as the timescale parameter to transform the continuous parameters $\mathbf{A}$ and $\mathbf{B}$ into discrete ones.
However, the constant dynamics in S4 prevent it from correctly selecting the context information. By introducing a selective mechanism, selective structured state space models (S6), also called Mamba mambagu2023mamba have been proposed to realize the data-dependent dynamics, thereby effectively filtering the context information and achieving a better performance in language and visual tasks.
2.3 Mamba in Vision
Motivated by Mamba’s excellent performance in language tasks, recent research has begun exploring its application in vision tasks and has achieved promising results. To address the unidirectional modeling and lack of positional awareness in the original Mamba model, Vision Mamba (Vim) vision_mamba incorporates the bidirectional SSMs for data-dependent global visual context modeling and position embeddings for location-aware visual recognition. VMamba liu2024vmamba introduces the Cross-Scan Module (CSM) to bridge the gap between 1-D array scanning and 2-D plain traversing. However, Vim and VMamba have not explored the role of SSM in the channel dimension.
More recently, some works have begun to apply the SSM mechanism to the channel dimension. CU-Mamba cumamba employs a Channel SSM component to preserve channel correlation features. S2Mamba s2mamba designs an extra scanning module on the spectral dimension for retrieving semantic properties in HSI data by a bi-directional interaction between each band. However, these methods directly conduct the scanning process along the spectral dimension and do not utilize the local similarity of HSIs in both spatial and spectral dimensions.
3 The Proposed ASLE-SSM Method
3.1 Architecture of ASLE-SSM
The overall architecture of our ASLE-SSM is depicted in Fig. 2. ASLE-SSM adopts a U-shaped network with the spatial-spectral SSM block as the basic unit. The structure consists of the encoder, bottleneck, and decoder. Given the vectorized measurement $\mathbf{y}\in\mathbb{R}^{H(W+d(N_{\lambda}-1))}$, we reverse the dispersion and shift back $\mathbf{Y}$ to initialize the input signal $\mathbf{X}\in\mathbb{R}^{H\times W\times N_{\lambda}}$:
$\mathbf{X}(x,y,n_{\lambda})=\mathbf{Y}(x,y-d(\lambda_{n}-\lambda_{c})).$ (6)
We first concatenate the input $\mathbf{X_{0}}$ with the physical mask $\mathbf{Y}\in\mathbb{R}^{H\times W\times N_{\lambda}}$ and then employ a 1$\times$1 convolution layer to integrate information from the original signal and the physical mask. Moreover, the integrated input $\mathbf{X_{0}}$ undergoes a 3$\times$3 convolution embedding into a feature $\mathbf{X_{0}^{{}^{\prime}}}\in\mathbb{R}^{H\times W\times C}$. Subsequently, $\mathbf{X_{0}^{{}^{\prime}}}$ passes through the downsample and upsample process. After passing through the decoder, we employ a 3$\times$3 convolution mapping to obtain the residual information. By computing the sum of the integrated input and residual output, we finally obtain the reconstructed HSIs. In our implementation, we change the number of spatial-spectral SSM blocks in the encoder, bottleneck, and decoder to realize the small, middle, and large sizes of ASLE-SSM models.
[…]
     Algorithms   Params   GFLOPs      S1      S2      S3      S4      S5      S6      S7      S8      S9      S10     Avg
GAP-TV gap_tv
26.04
0.817
21.66
0.724
24.86
0.732
30.51
0.875
24.33
0.778
25.11
0.790
18.28
0.730
23.94
0.780
21.77
0.732
23.08
0.721
23.96
0.768
DeSCI desci
28.38
0.803
26.00
0.701
23.11
0.730
28.26
0.855
25.41
0.778
24.66
0.764
24.96
0.725
24.15
0.747
23.56
0.701
24.17
0.677
25.27
0.748
$\lambda$-Net lambda
62.64M
117.98
30.10
0.849
28.49
0.805
27.73
0.870
37.01
0.934
26.19
0.817
28.64
0.853
26.47
0.806
26.09
0.831
27.50
0.826
27.13
0.816
28.53
0.841
TSA-Net tsa_net
44.25M
110.06
32.31
0.894
31.03
0.863
32.15
0.916
37.95
0.958
29.47
0.884
31.06
0.902
30.02
0.880
29.22
0.886
31.14
0.909
29.18
0.861
31.35
0.895
HDNet hdnet
2.37M
154.76
34.96
0.937
35.64
0.943
35.55
0.946
41.64
0.976
32.56
0.948
34.33
0.954
33.27
0.928
32.26
0.945
34.17
0.944
32.22
0.940
34.66
0.946
DIP-HSI self
33.85M
64.42
31.32
0.855
25.89
0.699
29.91
0.839
38.69
0.926
27.45
0.796
29.53
0.824
27.46
0.700
27.69
0.802
33.46
0.863
26.10
0.733
29.75
0.803
DiffSCI diffsci
93.52M
193.76
34.96
0.907
34.60
0.905
39.83
0.949
42.65
0.951
35.21
0.946
33.12
0.917
36.29
0.944
30.42
0.887
37.27
0.931
28.49
0.821
35.28
0.916
ADMM-Net admm-net
4.27M
78.58
34.03
0.919
33.57
0.904
34.82
0.933
39.46
0.971
31.83
0.924
32.47
0.926
32.01
0.898
30.49
0.907
33.38
0.917
30.55
0.899
33.26
0.920
DGSMP gsm
3.76M
646.65
33.26
0.915
32.09
0.898
33.06
0.925
40.54
0.964
28.86
0.882
33.08
0.937
30.74
0.886
31.55
0.923
31.66
0.911
31.44
0.925
32.63
0.917
GAP-Net gapnet
4.27M
78.58
33.63
0.913
33.19
0.902
33.96
0.931
39.14
0.971
31.44
0.921
32.29
0.927
31.79
0.903
30.25
0.907
33.06
0.916
30.14
0.898
32.89
0.919
MST-S mst
0.93M
12.96
34.67
0.936
34.46
0.927
34.75
0.940
41.54
0.972
32.10
0.937
33.62
0.945
32.75
0.918
31.77
0.938
34.00
0.937
31.74
0.930
34.14
0.938
MST-M mst
1.50M
18.07
35.05
0.942
35.25
0.943
35.75
0.955
42.27
0.979
32.49
0.948
34.28
0.955
33.09
0.923
32.49
0.951
34.82
0.950
32.30
0.944
34.78
0.949
MST-L mst
2.03M
28.15
35.28
0.946
35.93
0.948
36.19
0.955
41.64
0.977
32.83
0.950
34.61
0.957
33.95
0.932
32.80
0.953
34.86
0.947
32.66
0.946
35.07
0.951
CST-S cst
1.20M
11.67
34.57
0.935
34.61
0.934
35.26
0.949
41.42
0.975
32.14
0.942
34.43
0.955
33.31
0.925
32.75
0.951
34.79
0.949
32.06
0.939
34.53
0.945
CST-M cst
1.36M
16.91
34.96
0.943
35.48
0.945
36.30
0.957
41.99
0.979
32.74
0.951
35.11
0.961
33.70
0.930
33.39
0.958
35.14
0.952
32.71
0.947
35.15
0.952
CST-L cst
3.00M
27.81
35.61
0.951
36.32
0.955
37.20
0.963
42.09
0.980
33.28
0.956
35.48
0.967
34.25
0.940
33.62
0.965
35.74
0.957
33.26
0.955
35.69
0.959
CST-L+ cst
3.00M
40.10
35.64
0.951
36.79
0.957
37.71
0.965
41.38
0.981
32.95
0.957
35.58
0.966
34.54
0.947
34.07
0.964
35.62
0.959
32.82
0.949
35.71
0.960
ASLE-SSM-S
0.81M
7.01
35.01
0.938
34.57
0.932
36.19
0.953
42.04
0.976
32.40
0.943
33.99
0.949
33.42
0.928
32.38
0.945
35.54
0.951
31.74
0.933
34.74
0.946
ASLE-SSM-M
1.44M
11.14
35.33
0.945
35.45
0.945
36.99
0.962
43.19
0.982
32.44
0.949
34.49
0.958
33.65
0.936
33.16
0.957
35.21
0.954
32.32
0.943
35.22
0.953
ASLE-SSM-L
2.81M
14.51
35.66
0.951
36.03
0.952
37.92
0.967
43.44
0.983
33.03
0.956
34.92
0.963
34.26
0.940
33.47
0.961
36.08
0.962
32.56
0.950
35.74
0.958
3.2 Local Enhancement SSM (LE-SSM) Mechanism
Although Mamba is effective in modeling long-range dependencies, directly using the SSM mechanism often leads to less effective in extracting local features. Recent models tend to incorporate CNNs into the network structure, utilizing their local feature extraction capabilities to improve the performance of original SSMs. However, such methods would increase the parameter count, diminishing the lightweight advantage of the original SSM mechanism. While directly adding convolutional layers can enhance the model’s performance, it does not fundamentally solve the problem.
Inspired by LocalMamba localmamba , we introduce a simplified local mechanism in the spectral dimension. We use two scanning directions as the global branch, scanning the image row by row in both forward and backward directions while using the other two scanning directions as local branches to unfold and scan local patches. The specific details of the scanning process are illustrated in Fig. 2(d). Given the input $\mathbf{X}\in\mathbb{R}^{H\times W\times C}$, we first generate four copies of input data with different arrangements. For the global branch, we directly flatten the feature map into $\mathbf{X_{1}}\in\mathbb{R}^{L\times C}$, where $L=H\times W$, and then reverse it to generate $\mathbf{X_{2}}$. For the local branch, we first set the size of our patch flattening each patch into a 1D sequence, and then linking each patch together to generate $\mathbf{X_{3}}$. Similarly, we reverse $\mathbf{X_{3}}$ to obtain $\mathbf{X_{4}}$. After the generation of input data in four directions, we conduct the process as follows:
$\mathbf{X^{{}^{\prime}}}=Concat(\mathbf{X_{1}},\mathbf{X_{2}},\mathbf{X_{3}},\mathbf{X_{4}}),$ (7)
$\mathbf{\hat{X}}=LE\text{-}Spatial\text{-}SSM(\mathbf{X^{{}^{\prime}}}).$
The structure of LE-Spatial-SSM is shown in Fig. 2(b). It is worth noting that our method does not employ any CNNs or spatial and channel attention mechanisms based on MLPs, ensuring that our method does not add any extra parameters or time consumption compared to the global cross-scan mechanism liu2024vmamba . However, it also means we need to carefully choose the size of local patches, or else the reconstruction result will even become worse. In our implementation, we use patches of size 4$\times$4 to achieve a relatively better performance.
3.3 Across-Scanning SSM (AS-SSM) Mechanism
Our LE-SSM mechanism establishes a local-enhanced global receptive field but does not fully exploit the rich spectral information inherent in HSIs, leading to suboptimal reconstruction. Recently, some methods cumamba ; s2mamba have utilized and integrated channel information by directly scanning along the channel dimension (spectral dimension), but this scanning method does not consider the local features in both channel and spatial dimensions. In other words, neighboring pixels and bands often have certain similarities, but scanning the spectral dimension directly may lead to a distance increase between data points that are highly correlated.
To address the above issues, we propose a local scanning method that integrates local features from both spatial and spectral dimensions. This mechanism attempts to exploit the local properties of HSIs in both spatial and spectral dimensions while leveraging the rich spectral information to guide the reconstruction process. First, we rearrange the input $\mathbf{X}\in\mathbb{R}^{H\times W\times C}$ into patches along the spatial context by dividing the input HSIs into a series of smaller data cubes $\mathbf{S}\in\mathbb{R}^{P\times P\times C}$, which will be flattened into sequences for scanning. After that, we organize the scanning order within each sequence based on our spatial-spectral local cubes $\mathbf{C}\in\mathbb{R}^{h\times w\times c}$. As shown in Fig. 2(e), different colors indicate each spatial-spectral local cube. The scanning order of each data point is also depicted. The process can be formulated as:
$\mathbf{C}=Divide(\mathbf{X}),$ (8)
$\mathbf{C^{{}^{\prime}}}=AS\text{-}Spectral\text{-}Spatial\text{-}SSM(\mathbf{C}),$
where $\mathbf{C^{{}^{\prime}}}$ denotes the output patch of AS-SSM. After the SSM mechanism, we reorganize the patches back into the same order as the input $\mathbf{X}$. By scanning the data points within each data cube, we can leverage the local similarity between adjacent spectral bands and pixels, thus achieving better results than previous methods that directly scan along the spectral dimension.
4 Experiment
4.1 Experimental Settings
The effectiveness of our ASLE-SSM network is verified on both simulation and real datasets, by using 28 wavelengths ranging from 450nm to 650nm, acquired via spectral interpolation manipulation.
Simulation HSI Data. Two simulated HSI datasets CAVE cave and KAIST kaist are utilized to conduct the simulation experiment. The CAVE dataset contains 32 HSIs with a spatial size of $512\times 512$, while the KAIST dataset includes 30 HSIs with a spatial size of $2704\times 3376$. Following the experimental settings of TSA-Net tsa_net , we designate CAVE as the training set and select 10 scenes from the KAIST as the testing set.
Real HSI Data. We adopt the real HSIs obtained by the CASSI system, as developed in TSA-Net tsa_net for testing.
Evaluation Metrics. We select peak signal-to-noise ratio (PSNR) PSNR and structural similarity (SSIM) ssim as the evaluation metrics to assess the quality of HSI reconstruction.
Implementation Details. Our ASLE-SSM is implemented using PyTorch and we conduct model training with Adam optimizer ($\beta_{1}$ = 0.9 and $\beta_{2}$ = 0.999) for 300 epochs. The learning rate is set to $4\times 10^{-4}$ at the beginning and is halved every 50 epochs during the training procedure. Image patches with size $256\times 256$ are randomly cropped from the 3D HSI cubes with 28 channels for training in the simulation and real experiments. The shifting step $d$ in the dispersion process is set to 2 and the batch size is 2. We apply the data augmentation techniques including random flipping and rotation. The models are trained on an RTX 4090 GPU.
Refer to caption Scene 1 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Meas GT Patch ADMM-Net CST-L+ DGSMP DIP-HSI HDNet Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption GAP-Net DiffSCI $\lambda$-Net MST-L PnP-CASSI TSA-Net ASLE-SSM-L
Refer to caption Scene 2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Meas GT Patch ADMM-Net CST-L+ DGSMP DIP-HSI HDNet Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption GAP-Net DiffSCI $\lambda$-Net MST-L PnP-CASSI TSA-Net ASLE-SSM-L
Figure 4: Visual comparison on $\operatorname{KAIST}$. Top is $Scene$ 1 at wavelength 487.0nm. Bottom is $Scene$ 2 at wavelength 575.5nm
4.2 Simulation Results
Table 2: Comparisons of different frameworks on PSNR, SSIM, FLOPs, Params, and Inference Time on the test dataset.
Framework  MST-S mst  CST-S cst  ASLE-SSM-S
PSNR 34.14 34.53 34.74
SSIM 0.938 0.945 0.946
Params (M) 0.93 1.20 0.81
FLOPS (G) 12.96 11.67 7.01
Inference Time(s) 0.022 0.041 0.009
We conduct a comprehensive comparative analysis of our ASLE-SSM method and SOTA HSI reconstruction algorithms in Params, FLOPs, PSNR, and SSIM. We compared the results of ASLE-SSM with 12 SOTA methods, including two model-based methods (GAP-TV gap_tv , and DeSCI desci ), two CNN-based methods ($\lambda$-net lambda , TSA-Net tsa_net ), three Transformer-based methods (HDNet hdnet , MST-S, MST-M, MST-L mst , and CST-S, CST-M, CST-L, CST-L+ cst ), two Plug-and-Play (PnP) methods (DIP-HSI self , DiffSCI diffsci ) and three deep unfolding methods (ADMM-Net admm-net , GAP-Net gapnet , DGSMP gsm ) across 10 simulation scenes. The quantitative results of different models are listed in Table 1.
As shown in Table 1, our ASLE-SSM-L obtains 35.74dB of PSNR and 0.958 of SSIM, achieving the best reconstruction performance among SOTAs. When compared to CNN-based and Transformer-based methods such as TSA-net tsa_net , HDNet hdnet , MST-L mst , and CST-L+ cst , our ASLE-SSM-L demonstrates notable improvements over these methods, which are 4.39dB, 1.08dB, 0.67dB, and 0.03dB on average, respectively. The spectral density curves in Fig. 5 also demonstrate our superiority.
We also conduct comparisons of our ASLE-SSM-S with previous Transformer-based models MST mst and CST cst in the small size. As shown in Table 2, we not only achieve better performance in PSNR and SSIM compared to MST and CST but also obtain a more lightweight model with faster speed. Specifically, the ASLE-SSM-S is 2.4 times faster than MST-S and 4.5 times faster than CST-S, while having 0.12 (M) fewer parameters than the MST-S, which means the smallest parameter count. From Fig. 1, we can also observe that our ASLE-SSM not only achieves better results, but the value of FLOPs is also significantly lower than that of the corresponding MST and CST models.
Except for quantitative comparisons, we also visualize the results of simulation experiments. We choose one of 28 channels from Scene 1 and Scene 2, against the simulation results obtained from twelve SOTA approaches. Visually, as illustrated in Fig. 4, our method yields smoother and cleaner details. The experimental results indicate the capability of our method to generate high-quality HSIs. Specifically, in Scene 1, our reconstruction results better generate the curvature of the chin compared to other models. In Scene 2, our model clearly reconstructs the shapes of the blocks and ropes. The heart-like pattern on the blocks is also well-preserved.
4.3 Real Data Results
Figure 5: Spectral Density Curves.
We further apply our method to reconstruct real HSIs. We retrain our model on the CAVE cave and KAIST kaist datasets, conducting tests on real scenes for experiments. Following the configurations established in MST mst , We introduce 11-bit shot noise during training to simulate real measurement conditions. As shown in Fig. 6, we compare the reconstructed images from two real scenes (Scene 3 and Scene 4) using our ASLE-SSM and nine SOTA methods. Our model achieved superior results when compared to other models. Specifically, In Scene 3, our image preserves the most information compared to the results of other models. A noticeable protrusion consistent with the original image can be observed in the center of our result and other methods fail to reproduce.
Additionally, our images do not exhibit significant artifacts or ghosting. In Scene 4, We can observe that the face shape in the original image is relatively rounded. However, the reconstruction results of previous methods generate an almond-shaped face similar to the compressed measurements, indicating that these methods did not learn this information. In contrast, our results clearly show a fuller face on both sides. Additionally, previous reconstruction methods often only reconstruct one side of the mouth, while our image noticeably reconstructs the whole mouth, though there is a slight misalignment between the two sides. Furthermore, our images do not exhibit nonexistent features that appear in previous methods’ results, such as noses, wrinkles, or gaps that are not present in the original RGB image. These results demonstrate that our ASLE-SSM surpasses other methods in reconstructing image details and suppressing noise in real scenes.
[Uncaptioned image] Real Measurement [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] RGB Ref. Meas CST-L DGSMP GAP-Net HDNet [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] HLRTF DiffSCI $\lambda$-Net MST-L TSA-Net ASLE-SSM-L
Refer to caption Real Measurement Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption RGB Ref. Meas CST-L DGSMP GAP-Net HDNet Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption HLRTF DiffSCI $\lambda$-Net MST-L TSA-Net ASLE-SSM-L
Figure 6: Visual comparison of SCI reconstruction methods on real $Scene$ 3 at wavelength 648.0nm.
4.4 Ablation Study
Break-down Ablation. We conduct an ablation study by removing the LE-Spatial Mechanism and AS-Spectral-Spatial Mechanism from the ASLE-SSM network. The experiment results are shown in Table 3. Our goal is to explore the contributions of each mechanism to the results. The baseline barely adopts the U-shaped framework without any SSM modules.
Table 3: Ablation study of ASLE-SSM’s components
Architecture Params (M) PSNR SSIM
Baseline w/o SSM 0.18 27.41 0.678
with LE-SSM 0.62 34.48 0.942
with AS-SSM 0.37 31.16 0.863
with LE + AS SSM 0.81 34.74 0.946
When we respectively employ LE-SSM and AS-SSM blocks, our model achieves 7.07 dB and 3.75 dB improvements in PSNR over the baseline, respectively. Employing both mechanisms simultaneously can further enhance the model’s performance. Here we need to point out that, in consideration of maintaining lightweight, we only utilized a single-direction scan in AS-SSM. In fact, we could also deploy a four-direction scanning similar to LE-SSM. However, each scan direction implies additional computational costs so we decided not to adopt this way.
Table 4: Comparisons of AS mechanism and common scanning.
Architecture PSNR SSIM
w/o Spectral SSM 34.48 0.942
with AS Mechanism 34.74 0.946
with Common Scanning 34.55 0.943
Across-Scanning Mechanism. To demonstrate the advantages of our across-scanning mechanism compared to common spectral scanning, we conduct the experiment on model performance under both methods. As illustrated in Table 4, when we adopt common scanning, it yields a 0.07 dB improvement. In contrast, when we apply the proposed across-scanning mechanism, our method achieves a significant increase of 0.26 dB compared to not using spectral SSM. Both methods were experimented with under the condition of a single scan. The superiority of our method mainly comes from paying more attention to adjacent bands and pixels, thus better extracting the local features of HSIs along both spatial and spectral dimensions. The results verify the role of our AS mechanism.
Table 5: Comparisons of LE mechanism and global scanning.
Architecture PSNR SSIM
w/o Spatial SSM 31.16 0.863
with LE Machenism 34.48 0.942
with Global Scanning 34.40 0.940
Local-Enhancement Mechanism. Additionally, we conduct an ablation study of our LE mechanism on the CAVE and KAIST datasets to assess its contributions. As shown in Table 5, The evaluations begin with the model without SSM in the spatial dimension, which served as our initial baseline with a PSNR of 31.16 dB. Specifically, the addition of the global four-direction scanning along the spatial context yields a significant increase of 3.32 dB in PSNR and 0.077 in SSIM, demonstrating its effectiveness in modeling the global spatial context.
We further apply our LE Machenism along the spatial dimension and achieve a 0.08 dB increase compared to the global scanning. our method only changes the scanning order and does not introduce any additional parameter consumption or computational cost to achieve a relatively better performance. In other words, we do not use CNN layers or spatial attention mechanisms as in other methods localmamba . In fact, Without increasing any additional cost and with the reconstructed results already achieving a relatively good level, 0.08 dB in PSNR and 0.002 in SSIM are significant improvements. The ablation experiment demonstrates the importance of local features in HSIs and verifies the effectiveness of our LE Mechanism.
5 Conclusion
In this paper, we address two limitations of previous SSM-based models by our proposed mechanisms: insufficient utilization of channel (spectral) information, and limited local detail extraction capability. By designing a novel across-scanning mechanism, we fully utilize the rich spectral information in HSIs to guide the image reconstruction process. At the same time, our scanning method based on patched data cubes pays more attention to the local similarity of HSIs between adjacent bands and pixels, achieving better results compared to simply scanning along the channel dimension. By introducing local enhancement in the Spatial SSM, we integrate spatial context from both global and local perspectives, achieving better results than using a global SSM alone.
Despite the significant progress of our method, we also acknowledge that as the first framework to adopt the SSM mechanism in HSI reconstruction, our network still has room for further improvement. For example, we could design a lightweight and effective way to fuse local and global branch information in the spatial dimension, rather than simply adding them together. However, the key challenge lies not in improving the effect itself, but in controlling computational costs. We will explore this aspect in the future with the aim of designing a more efficient network architecture.