### v2

# Cross-Scan Mamba with Masked Training for Robust Spectral Imaging

Abstract
Snapshot Compressive Imaging (SCI) enables fast spectral imaging but requires effective decoding algorithms for hyperspectral image (HSI) reconstruction from compressed measurements. Current CNN-based methods are limited in modeling long-range dependencies, while Transformer-based models face high computational complexity. Although recent Mamba models outperform CNNs and Transformers in RGB tasks concerning computational efficiency or accuracy, they are not specifically optimized to fully leverage the local spatial and spectral correlations inherent in HSIs. To address this, we propose the Cross-Scanning Mamba, named CS-Mamba, that employs a Spatial-Spectral SSM for global-local balanced context encoding and cross-channel interaction promotion. Besides, while current reconstruction algorithms perform increasingly well in simulation scenarios, they exhibit suboptimal performance on real data due to limited generalization capability. During the training process, the model may not capture the inherent features of the images but rather learn the parameters to mitigate specific noise and loss, which may lead to a decline in reconstruction quality when faced with real scenes. To overcome this challenge, we propose a masked training method to enhance the generalization ability of models. Experiment results show that our CS-Mamba achieves state-of-the-art performance and the masked training method can better reconstruct smooth features to improve the visual quality.
1 Introduction
Figure 1: PSNR-FLOPs comparisons of CS-Mamba and previous Deep Unfolding SOTA methods.
Hyperspectral images (HSIs) contain a large number of spectral bands to store rich detail information and thus are widely used in various fields such as object tracking [25, 36, 40], remote sensing [9, 21, 49], super-resolution [12, 44, 20], and medical image processing [1, 30, 33]. Traditional imaging systems used for collecting HSIs involve scanning along the spatial or spectral dimensions but this imaging process is time-consuming and can not effectively capture dynamic objects. Recently, snapshot compressive imaging (SCI) systems have been developed [29, 41] as an alternative method for obtaining HSIs at video rate. SCI systems capture multiple image bands in a single snapshot, compressing the information of HSIs into a 2D measurement. [42, 51, 6, 15]. Among these systems, Coded Aperture Snapshot Spectral Imaging (CASSI) has emerged and become mainstream [50, 48]. CASSI utilizes a physical mask to modulate signals from HSIs, shifting these modulated images to different spatial locations on the detector plane via a disperser [3, 47]. However, the performance of the CASSI system heavily depends on the efficiency of the reconstruction algorithm.
In recent years, numerous reconstruction algorithms have been proposed, including model-based methods and deep learning-based methods. Traditional model-based methods typically attempt to model the inverse process of image reconstruction, using manually designed priors to derive the solution to an optimization problem. Although these methods are highly interpretable, they often require manual parameter tuning and tend to be slow. With the success of deep learning methods in various visual tasks, CNN and Transformer-based methods have been developed and applied to HSI reconstruction, achieving higher image reconstruction quality and speed than traditional algorithms. The CNN-based methods [35, 32] have an advantage in extracting local features from images, but they lack the ability to model long-range dependencies effectively. Transformer-based methods [3, 4, 5, 2] realize a global receptive field, resulting in superior performance compared to CNNs but higher computational costs. Although some mechanisms can mitigate this issue, they do not fundamentally change the quadratic complexity problem caused by the attention mechanism in the Transformer architecture.
The limitation of Transformers in spectral SCI remains unresolved. However, in the realm of large language model and RGB image vision, an advanced solution for this problem exists, known as Mamba [16]. Through data-dependent selective SSMs with parallel scanning and hardware-wise strategies, Mamba establishes a global receptive field and achieves near-linear complexity. Mamba-based methods have been shown to outperform previous methods in various tasks due to their excellent long sequence modeling capabilities and lower computational costs [45, 54, 19, 27].
Subsequently, many works have begun to explore the application and performance of the Mamba model in various visual tasks. Representative works, such as the visual state space model [39] and vision Mamba [53], have demonstrated Mamba’s superiority for image processing, particularly in obtaining fewer parameters and faster speed. However, most Mamba-based models only conduct the scanning process along the spatial dimension, lacking cross-channel interaction, despite the importance of spectral information for HSI reconstruction. Recently, some Mamba-based methods have made progress in developing an SSM mechanism along the channel (spectral) dimension for certain visual tasks [43, 22], achieving better results compared to scanning only along the spatial dimension. However, this scanning method does not consider the inherently local features of HSIs in both spatial and spectral dimensions. Additionally, previous reconstruction algorithms have achieved better results in experiments by continuously improving the backbone and framework. However, it is often overlooked that achieving superior performance in the simulation experiment is not entirely linearly correlated with achieving better visual effects in the real experiment. Simulation experiments primarily address the losses caused by the reconstruction compression process, whereas real experiments introduce noise that simulates real-world conditions, thus requiring the model to possess stronger generalization capabilities.
However, the degradation caused by insufficient generalization capability is, in fact, not easy to notice. As a result, most current work has not recognized this issue. This oversight is mainly due to the limited test data from real-world experiments, which makes certain issues less likely to be noticed. In most scenes, an improvement in quantitative results in simulation experiments is indeed associated with better performance in real scenes, which has led to some confusion. In this paper, we provide a detailed explanation of this phenomenon, making the problem more apparent. Notably, certain optimization strategies commonly used in previous simulation experiments can even deteriorate specific features in real-world scenes. We assume this phenomenon arises due to the model’s overemphasizing of compression loss learning, thereby neglecting other noise (like shot noise) and the semantic information inherent in images. Experimental results in this paper confirm this point.
In this paper, we propose CS-Mamba, a spatial-spectral SSM with our cross-scanning method and local enhancement, to improve HSI reconstruction in SCI. Our CS-Mamba aims to strike a balance between global and local information utilization in HSIs. CS-Mamba comprises two main components: spatial SSM and spectral-spatial cross SSM. The spatial SSM employs a combination of global and local scanning methods, enhancing row-by-row scanning with both global and local modeling. Meanwhile, the spectral-spatial cross SSM incorporates an innovative across-scanning mechanism, leveraging spatial-spectral local cubes to enhance the module’s ability to capture local details. Besides, we introduce a masked training strategy in real scenarios. This strategy forces the model to learn the inherent features of images, enhancing the generalization capabilities of previous methods and thereby achieving better results in real scenes. Overall, our contributions are summarized as follows:
  • We present a novel spectral-spatial state space model for HSI reconstruction, which can simultaneously integrate information from both spectral and spatial dimensions to adequately model spatial context and spectral features.
  • We propose a cross-scanning method that integrates local features from both spatial and spectral dimensions within a single scan. Additionally, we introduce a local scanning approach within the spatial SSM to further enhance local feature extraction.
  • We introduce a masked training strategy to promote the robustness of models in real scenarios.
Figure 2: Schematic of the CASSI system: A 3D HSI cube is modulated by a mask, sheared by a disperser, and transformed into a 2D-coded measurement.
2 Preliminaries and Related Works
CASSI System. An imaging process CASSI system is illustrated in Fig. 2. The 3D HSI data cube is denoted as $\mathbf{F}\in\mathbb{R}^{H\times W\times N_{\lambda}}$ where H, W, and $N_{\lambda}$ represent the height, width, and number of wavelengths of the 3D HSI, respectively. A physical mask $\mathbf{M}\in\mathbb{R}^{H\times W}$ is used to modulate the 3D HSI $\mathbf{F}$ as: $\mathbf{F^{{}^{\prime}}}(:,:,n_{\lambda})=\mathbf{F^{{}^{\prime}}}(:,:,n_{\lambda})\odot\mathbf{M},$ where $\odot$ represents the element-wise product and $\mathbf{F^{{}^{\prime}}}\in\mathbb{R}^{H\times W\times N_{\lambda}}$ is the modulated 3D HSIs. After the modulated cube passes the disperser, $\mathbf{F^{{}^{\prime}}}$ is tilted and considered to be sheared along the y-axis. We use $\mathbf{F^{{}^{\prime\prime}}}\in\mathbb{R}^{H\times(W+d(N_{\lambda}-1))\times N_{\lambda}}$ to denote the tilted cube and assume $n_{\lambda}$ to be the reference wavelength. That is, image $\mathbf{F^{{}^{\prime}}}(:,:,n_{\lambda_{c}})$ is not sheared along the y-axis. Then we have: $\mathbf{F^{{}^{\prime\prime}}}(u,v,n_{\lambda})=\mathbf{F^{{}^{\prime}}}(x,y+d(\lambda_{n}-\lambda_{c}),n_{\lambda})$, where $(u,v)$ indicates the coordinate system on the detector plane, and $\lambda_{n}$ is the wavelength of channel $n_{\lambda}$. Here, $d(\lambda_{n}-\lambda_{c})$ signifies the spatial shifting for band $n_{\lambda}$. Finally, the measurement $\mathbf{Y}\in\mathbb{R}^{H\times[W+d(N_{\lambda}-1)]}$ can be obtained by
$\mathbf{Y}=\sum_{n_{\lambda}=1}^{N_{\lambda}}\mathbf{F^{{}^{\prime\prime}}}(:,:,n_{\lambda})+\mathbf{N},$ (1)
where $\mathbf{N}\in\mathbb{R}^{H\times[W+d(N_{\lambda}-1)]}$ represents the noise of CASSI caused by the sensing detector.
State Space Models. Recent advancements in state space models (SSMs) have spurred the development of various innovative methods and architectures. Notably, the structured state space model (S4) [18] has emerged as a pioneering approach due to its effectiveness in capturing long-range dependencies. S4 maps a 1D sequence $x(t)\in\mathbb{R}\rightarrow y(t)\in\mathbb{R}$ by an implicit latent state $h(t)\in\mathbb{R}^{N}$ and decodes it into an output sequence, which can be formulated as:
$h^{\prime}(t)=\mathbf{A}h(t)+\mathbf{B}x(t),$ (2)
$y(t)=\mathbf{C}h(t)+\mathbf{D}x(t),$
where $N$ is the state size, ${\rm\textbf{A}}\in\mathbb{R}^{N\times N}$, ${\rm\textbf{B}}\in\mathbb{R}^{N\times 1}$, ${\rm\textbf{C}}\in\mathbb{R}^{1\times N}$, and ${\rm\textbf{D}}\in\mathbb{R}$. After that, a discretization process is adopted to integrate Eq. (2) into the deep learning algorithm, using the zero-order hold (ZOH) rule defined as,
$\overline{\rm\textbf{A}}$ $={\rm exp}({\rm{\Delta\textbf{A}}}),$ (3)
$\overline{\rm\textbf{B}}$ $=({\rm{\Delta\textbf{A}}})^{-1}({\rm exp(\textbf{A})}-\textbf{I})\cdot{\rm\Delta\textbf{B}},$
where $\overline{\rm\textbf{A}}$ and $\overline{\mathbf{B}}$ represent the discretized forms of the parameters $\mathbf{A}$ and $\mathbf{B}$, respectively. $\rm\Delta$ is denoted as the timescale parameter to transform the continuous parameters $\mathbf{A}$ and $\mathbf{B}$ into discrete ones.
Mamba in Vision. Selective structured state space models (S6), or Mamba [17], use a selective mechanism for data-dependent dynamics, enhancing language and visual task performance. Vision Mamba (Vim) [53] addresses Mamba’s unidirectional modeling by adding bidirectional SSMs for global context and position embeddings for visual recognition. VMamba [28] bridges 1-D and 2-D scanning via the Cross-Scan Module (CSM). CU-Mamba [11] adds Channel SSM to capture channel correlations, while S2Mamba [43] introduces spectral scanning for semantic retrieval in HSI data. However, these methods overlook local similarity in both spatial and spectral dimensions of HSIs.
Figure 3: The flowchart of the proposed CS-Mamba. (a) Parameter Estimation Network. This module is used for capturing parameter $\alpha$ and noise level $\beta$ to guide the reconstruction process. (b) The U-shaped denoiser network and DUN framework. (c) The Spatial-Spectral SSM Block adopted in (b). This module mainly consists of a spatial SSM and a spectral SSM, depicted in the sub-flowchart (d) and (e), respectively. (f) The specific method of LE-SSM in the Spatial SSM Module. Scanning directions are divided into global and local branches. (g) The illustration of our Cross-Scan Mechanism. The scanning process is demonstrated by labeling the order of each data point.
3 The Proposed CS-Mamba Method
3.1 Architecture of CS-Mamba
The architecture of our CS-Mamba is illustrated in Fig. 3. CS-Mamba leverages a deep unfolding framework guided by mask degradation patterns. Using MAP theory, we model the HSI reconstruction process as follows:
$\mathbf{\hat{x}}=\text{arg}~{}\underset{\mathbf{x}}{\text{min}}~{}~{}\frac{1}{2}||\mathbf{y}-\mathbf{\Phi}\mathbf{x}||^{2}+\tau R(\mathbf{x}),$ (4)
where $||\mathbf{y}-\mathbf{\Phi}\mathbf{x}||^{2}$ is the data fidelity term, $R(\mathbf{x})$ represents the image prior, and $\tau$ controls the balance between the data fidelity and prior terms.
HQS introduces an auxiliary variable $z$, resulting in a constrained optimization problem given by
$\hat{x}=\arg\min_{x}\frac{1}{2}\|y-\Phi x\|^{2}+\lambda R(z)\quad s.t.\quad z=x.$ (5)
Eq. (5) is solved by minimizing
$\mathcal{L}_{\mu}(x,z)=\frac{1}{2}\|y-\Phi x\|^{2}+\lambda R(z)+\frac{\mu}{2}\|z-x\|^{2},$ (6)
where $\mu$ is a penalty parameter. Subsequently, Eq. (6) can be addressed by iteratively solving a data subproblem (x-subproblem) and a prior subproblem (z-subproblem):
${x_{k}=\arg\min_{x}\|y-\Phi x\|^{2}+\mu\|x-z_{k-1}\|^{2},}$ (7a)
${z_{k}=\arg\min_{z}\frac{1}{2(\sqrt{\lambda/\mu})^{2}}}\|z-x_{k}\|^{2}+R(z).$ (7b)
The data subproblem, Eq. (7a), usually has a fast closed-form solution as
$x_{k}=(\Phi^{\top}\Phi+\mu I)^{-1}(\Phi^{\top}y+\mu z_{k-1}),$ (8)
where $I$ is an identity matrix, the matrix inversion in Eq. (8) can be written as
$(\Phi^{\top}\Phi+\mu I)^{-1}=\mu^{-1}I-\mu^{-1}\Phi^{\top}(I+\Phi\mu^{-1}\Phi^{\top})^{-1}\Phi\mu^{-1}.$ (9)
In CASSI systems, $\Phi\in\mathbb{R}^{HW^{\prime}\times HW^{\prime}N_{\lambda}}$ is a block diagonal matrix, thus, $\Phi\Phi^{\top}$ is a diagonal matrix that can be defined as
$\Phi\Phi^{\top}\stackrel{{\scriptstyle def}}{{=}}\text{diag}\{\phi_{1},...,\phi_{i},...,\phi_{HW^{\prime}}\}.$ (10)
By plugging $\Phi\Phi^{\top}$ into $(I+\Phi\mu^{-1}\Phi^{\top})^{-1}$, we obtain:
$(I+\Phi\mu^{-1}\Phi^{\top})^{-1}=\text{diag}\{\frac{\mu}{\mu+\phi_{1}},...,\frac{\mu}{\mu+\phi_{i}},...,\frac{\mu}{\mu+\phi_{HW^{\prime}}}\}.$ (11)
By plugging Eq. (9), Eq. (10) and Eq. (11) into Eq. (8) and simplifying the formula, we have
$x_{k}=z_{k-1}+\Phi^{\top}[(y-\Phi z_{k-1})./(\mu+diag(\Phi\Phi^{\top}))],$ (12)
where $diag$ represents the extraction of elements from its diagonal to form a vector.
The prior subproblem Eq. (7b), from a Bayesian perspective, corresponds to Gaussian denoising on $x_{k}$ with noise level $\sqrt{\lambda/\mu}$. To address this, Eq. (7b) could be rewritten as,
$z_{k}=Denoiser(x_{k},\sqrt{\lambda/\mu}).$ (13)
DUNs iteratively solve the data subproblem (Eq. 12) and the prior subproblem (Eq. 13) to address ill-posed inverse problems. The data subproblem depends on the degradation matrix, which can be estimated using either a sensing matrix [3, 2] or a DNN [5]. However, sensing matrices may not accurately reflect real-world degradations, and directly modeling them is challenging. The prior subproblem is treated as a denoising task, but the noise level is unknown in reconstruction, affecting performance. Following [14], we iteratively update the degradation matrix and parameter estimates, learning constant noise and regularization parameters towards each stage due to computational cost, as detailed in the supplementary material.
For the denoiser network, we employ a customized Mamba using a spatial-spectral SSM block as its fundamental unit. The architecture comprises an encoder, a bottleneck, and a decoder. Given a measurement $\mathbf{y}\in\mathbb{R}^{H(W+d(N_{\lambda}-1))}$, we reverse the dispersion and shift back $\mathbf{Y}$ to initialize the input signal $\mathbf{X}\in\mathbb{R}^{H\times W\times N_{\lambda}}$:
$\mathbf{X}(x,y,n_{\lambda})=\mathbf{Y}(x,y-d(\lambda_{n}-\lambda_{c})).$ (14)
We first concatenate the input $\mathbf{X_{0}}$ with the physical mask $\mathbf{Y}\in\mathbb{R}^{H\times W\times N_{\lambda}}$ and then employ a 1$\times$1 convolution layer to integrate information from the original signal and the physical mask. Moreover, the integrated input $\mathbf{X_{0}}$ undergoes a 3$\times$3 convolution embedding into a feature $\mathbf{X_{0}^{{}^{\prime}}}\in\mathbb{R}^{H\times W\times C}$. $\mathbf{X_{0}^{{}^{\prime}}}$ undergoes downsampling and upsampling before entering the decoder, followed by a 3$\times$3 convolution to extract residual information. Summing the integrated input and residual output yields the reconstructed HSIs. We adjust the number of spatial-spectral SSM blocks in the encoder, bottleneck, and decoder to create different CS-Mamba model sizes.
3.2 Enhanced Local SSM (LE-SSM)
Although Mamba effectively models long-range dependencies, directly using the SSM mechanism often limits its ability to capture local features. Recent methods integrate CNNs into the network to leverage their local feature extraction capabilities, enhancing original SSM performance. However, this increases the parameter count, undermining the lightweight nature of SSMs. While adding CNN can boost performance, it does not address the core issue. Inspired by LocalMamba [24], we propose a simplified local mechanism in the spectral dimension. This involves using two global scanning directions, traversing the image row by row in both forward and backward directions, while the other two scanning directions serve as local branches to unfold and scan local patches.
Figure 4: A visual comparison of with and w/o masked methods in our CS-Mamba model on real scene 1.
Table 1: Comparisons of our denoiser with previous End-to-End networks on PSNR, SSIM, FLOPs, Params, and Inference Time on the test dataset.
Framework  MST-S [3]  CST-S [2]  Cross-Scan SSM
PSNR 34.14 34.53 34.74
SSIM 0.938 0.945 0.946
Params (M) 0.93 1.20 0.81
FLOPS (G) 12.96 11.67 7.01
Inference Time(s) 0.022 0.041 0.009
The specific details of the scanning process are illustrated in Fig. 3(d). Given the input $\mathbf{X}\in\mathbb{R}^{H\times W\times C}$, we first generate four copies of input data with different arrangements. For the global branch, we directly flatten the feature map into $\mathbf{X_{1}}\in\mathbb{R}^{L\times C}$, where $L=H\times W$, and then reverse it to generate $\mathbf{X_{2}}$. For the local branch, we first set the size of our patch flattening each patch into a 1D sequence, and then linking each patch together to generate $\mathbf{X_{3}}$. Similarly, we reverse $\mathbf{X_{3}}$ to obtain $\mathbf{X_{4}}$. After the generation of input data in four directions, we conduct the process as follows:
$\mathbf{X^{{}^{\prime}}}=Concat(\mathbf{X_{1}},\mathbf{X_{2}},\mathbf{X_{3}},\mathbf{X_{4}}),$ (15)
$\mathbf{\hat{X}}=LE\text{-}SSM(\mathbf{X^{{}^{\prime}}}).$
The structure of LE-SSM is shown in Fig. 3(b). It is worth noting that our method does not employ any CNNs or spatial and channel attention mechanisms based on MLPs, ensuring that our method does not add any extra parameters or time consumption compared to the global cross-scan mechanism [28]. However, it also means we need to carefully choose the size of local patches, or else the reconstruction result will even become worse. In our implementation, we use patches of size 4$\times$4 to achieve a relatively better performance.
3.3 Cross-Scanning SSM (CS-SSM) Mechanism
Our LE-SSM mechanism creates a locally enhanced global receptive field but does not fully utilize the rich spectral information in HSIs. Recent methods [11, 43] have integrated channel information by scanning along the spectral dimension; however, this approach overlooks local features in both spectral and spatial dimensions. Neighboring pixels and bands often share similarities, but direct spectral scanning can increase the distance between highly correlated data points, reducing reconstruction quality.
Table 2: Comparisons of our CS-Mamba-S (9stg) with previous Transformer-based methods on PSNR, SSIM, FLOPs, and Params.
Framework  DAUHST [5]  DERNN [14]  CS-Mamba-S
PSNR 38.36 39.93 40.00
SSIM 0.967 0.976 0.980
Params (M) 6.15M 0.65M 0.57M
FLOPS (G) 79.50 81.99 65.77
To address the above issues, we propose a local scanning method that integrates local features from both spatial and spectral dimensions. This mechanism attempts to exploit the local properties of HSIs in both spatial and spectral dimensions while leveraging the rich spectral information to guide the reconstruction process. First, we rearrange the input $\mathbf{X}\in\mathbb{R}^{H\times W\times C}$ into patches along the spatial context by dividing the input HSIs into a series of smaller data cubes $\mathbf{S}\in\mathbb{R}^{P\times P\times C}$, which will be flattened into sequences for scanning. After that, we organize the scanning order within each sequence based on our spatial-spectral local cubes $\mathbf{C}\in\mathbb{R}^{h\times w\times c}$. As shown in Fig. 3(e), different colors indicate each spatial-spectral local cube. The scanning order of each data point is also depicted. The process can be formulated as follows:
$\mathbf{C}=Divide(\mathbf{X}),$ (16)
$\mathbf{C^{{}^{\prime}}}=CS\text{-}SSM(\mathbf{C}),$
where $\mathbf{C^{{}^{\prime}}}$ denotes the output patch of CS-SSM. After the SSM mechanism, we reorganize the patches back into the same order as the input $\mathbf{X}$. By scanning the data points within each data cube, we can leverage the local similarity between adjacent spectral bands and pixels, thus achieving better results than previous methods that directly scan along the spectral dimension.
[…]
     Algorithms   Reference   Params   GFLOPs      S1      S2      S3      S4      S5      S6      S7      S8      S9      S10     Avg
DIP-HSI [34] ICCV 2021
33.85M
64.42
31.32
0.855
25.89
0.699
29.91
0.839
38.69
0.926
27.45
0.796
29.53
0.824
27.46
0.700
27.69
0.802
33.46
0.863
26.10
0.733
29.75
0.803
ADMM-Net [31] ICCV 2019
4.27M
78.58
34.03
0.919
33.57
0.904
34.82
0.933
39.46
0.971
31.83
0.924
32.47
0.926
32.01
0.898
30.49
0.907
33.38
0.917
30.55
0.899
33.26
0.920
DGSMP [23] TPAMI 2023
3.76M
646.65
33.26
0.915
32.09
0.898
33.06
0.925
40.54
0.964
28.86
0.882
33.08
0.937
30.74
0.886
31.55
0.923
31.66
0.911
31.44
0.925
32.63
0.917
DiffSCI [37] CVPR 2024
93.52M
193.76
34.96
0.907
34.60
0.905
39.83
0.949
42.65
0.951
35.21
0.946
33.12
0.917
36.29
0.944
30.42
0.887
37.27
0.931
28.49
0.821
35.28
0.916
MST-L [3] CVPR 2022
2.03M
28.15
35.28
0.946
35.93
0.948
36.19
0.955
41.64
0.977
32.83
0.950
34.61
0.957
33.95
0.932
32.80
0.953
34.86
0.947
32.66
0.946
35.07
0.951
CST-L+ [2] ECCV 2022
3.00M
40.10
35.64
0.951
36.79
0.957
37.71
0.965
41.38
0.981
32.95
0.957
35.58
0.966
34.54
0.947
34.07
0.964
35.62
0.959
32.82
0.949
35.71
0.960
DAUHST-9stg [5] NIPS 2022 6.15M 79.50
37.25
0.958
39.02
0.967
41.05
0.971
46.15
0.983
35.80
0.969
37.08
0.970
37.57
0.963
35.10
0.966
40.02
0.970
34.59
0.956
38.36
0.967
PADUT [26] ICCV 2023 5.38M 90.46
37.36
0.962
40.43
0.978
42.38
0.979
46.62
0.990
36.26
0.974
37.27
0.974
37.83
0.966
35.33
0.974
40.86
0.978
34.55
0.963
38.89
0.974
RDLUF [13] CVPR 2023 1.89M 115.34
37.94
0.966
40.95
0.977
43.25
0.979
47.83
0.990
37.11
0.976
37.47
0.975
38.58
0.969
35.50
0.970
41.83
0.978
35.23
0.962
39.57
0.974
DERNN (9stg) [14] TGRS 2024 0.65M 81.99
38.26
0.965
40.97
0.979
43.22
0.979
48.10
0.991
38.08
0.980
37.41
0.975
38.83
0.971
36.41
0.973
42.87
0.981
35.15
0.962
39.93
0.976
DERNN (9stg)* [14] TGRS 2024 1.04M 134.18
38.49
0.968
41.27
0.980
43.97
0.980
48.61
0.992
38.29
0.981
37.81
0.977
39.30
0.973
36.51
0.974
43.38
0.983
35.61
0.966
40.33
0.977
SSR-L* [52] CVPR 2024 1.73M 78.93
38.81
0.968
41.51
0.979
43.76
0.979
48.62
0.988
38.32
0.979
37.85
0.975
38.50
0.969
36.85
0.974
42.64
0.980
35.82
0.965
40.27
0.976
CS-Mamba (3stg) Ours
0.97M
37.80
37.93
0.968
40.00
0.978
42.73
0.983
47.18
0.993
36.48
0.976
37.10
0.976
37.91
0.970
35.39
0.974
41.43
0.981
34.82
0.965
39.10
0.977
CS-Mamba (5stg) Ours
0.97M
62.93
38.23
0.971
40.62
0.982
43.44
0.984
47.58
0.994
37.95
0.982
37.74
0.979
38.71
0.974
36.27
0.979
42.89
0.986
35.45
0.970
39.90
0.980
CS-Mamba (9stg) Ours
0.97M
113.18
38.51
0.973
41.30
0.984
43.77
0.985
47.55
0.994
38.72
0.985
38.05
0.981
38.88
0.976
36.64
0.981
43.06
0.987
35.54
0.971
40.21
0.982
CS-Mamba* (9stg) Ours
1.16M
124.98
38.68
0.974
41.55
0.985
43.99
0.986
47.86
0.994
38.85
0.986
38.15
0.982
39.09
0.976
36.86
0.983
43.25
0.988
35.78
0.971
40.42
0.983
CS-Mamba-s (9stg) Ours
0.57M
65.77
38.63
0.972
40.90
0.983
43.46
0.984
47.30
0.994
38.41
0.983
37.68
0.979
38.80
0.975
36.29
0.979
43.09
0.986
35.41
0.969
40.00
0.980
Figure 5: A visual comparison of the performance of MST, CST, and DAUHST methods in the real scene.
3.4 Masked Training Strategy
Before presenting our masked training strategy, we discuss our observed phenomena and motivation for this approach. In Fig. 5, we display visual results for three Transformer-based methods [3, 2, 5] in a real scene. While their simulation performance improves from left to right, visual quality degrades in real scenarios. Compared to MST, CST’s reconstructed image shows more noise artifacts on the nose and forehead. These spots, absent in ground truth, suggest noise from 2D compressed measurements. DAUHST’s results exhibit even greater degradation, which is consistently observed across other algorithms in this scene.
This effect stems from two factors. First, simulation-based image reconstruction, aimed at minimizing compression loss, struggles with unknown real-world noise distributions. While shot noise is added during training, it inadequately reflects real-world complexities, leading to insufficient generalization. Secondly, focusing on mask pattern recovery can overemphasize specific information, weakening resilience to other noise types, and creating a balance challenge between compression and real-world losses.
$Total\ Loss=Compression\ Loss+\epsilon,$ (17)
where $\epsilon$ represents the complex real noise.
We propose a masked training strategy inspired by [8] for SCI reconstruction to enhance robustness and generalization. During training, a randomly generated 0-1 mask is applied to feature maps, creating an inpainting task that compels the network to reconstruct occluded regions. This strengthens the model’s capacity to handle complex real-world scenarios. Unlike traditional image denoising, where noise is added to clean images, SCI reconstruction starts with compressed 2D measurements that inherently carry compression loss, compounded by mixed noise in practice. By consistently applying the same mask during both training and testing, we ensure alignment between phases, leading to improved reconstruction accuracy and visually appealing results. This approach addresses unique SCI challenges, enhancing model performance under varied real-world conditions.
4 Experiment
4.1 Experimental Settings
The effectiveness of our CS-Mamba network is verified on both simulation and real datasets, by using 28 wavelengths ranging from 450nm to 650nm, acquired via spectral interpolation manipulation.
Simulation HSI Data. Two simulated HSI datasets CAVE [38] and KAIST [10] are utilized to conduct the simulation experiment. The CAVE dataset contains 32 HSIs with a spatial size of $512\times 512$, while the KAIST dataset includes 30 HSIs with a spatial size of $2704\times 3376$. Following the experimental settings of TSA-Net [32], we designate CAVE as the training set and select 10 scenes from the KAIST as the testing set.
Real HSI Data. We adopt the real HSIs obtained by the CASSI system, as developed in TSA-Net [32] for testing.
Evaluation Metrics. We select peak signal-to-noise ratio (PSNR) [7] and structural similarity (SSIM) [46] as the evaluation metrics to assess the quality of HSI reconstruction.
Implementation Details. Our CS-Mamba is implemented using PyTorch and we conduct model training with Adam optimizer ($\beta_{1}$ = 0.9 and $\beta_{2}$ = 0.999) for 300 epochs. The learning rate is set to $1\times 10^{-3}$ at the beginning and a cosine annealing scheduler is adopted. Image patches with size $256\times 256$ are randomly cropped from the 3D HSI cubes with 28 channels for training in the simulation and real experiments. The shifting step $d$ in the dispersion process is set to 2. We apply the data augmentation techniques including random flipping and rotation. Models are trained on RTX 4090 GPUs.
Refer to caption Scene 2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Meas GT Patch ADMM-Net CST-L+ DGSMP DIP-HSI HDNet Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption GAP-Net DiffSCI $\lambda$-Net MST-L DAUHST PADUT CS-Mamba
Figure 6: Visual comparison of different methods on simulation $Scene$ 2 at wavelength 575.5nm
4.2 Simulation Results
We conduct a comprehensive comparative analysis of our CS-Mamba method and SOTA HSI reconstruction algorithms in Params, FLOPs, PSNR, and SSIM. We compared the results of CS-Mamba with 11 SOTA methods algorithms across 10 simulation scenes. The quantitative results of different models are listed in Table 3.
As shown in Table 3, our CS-Mamba-9stg* obtains 40.42dB of PSNR and 0.983 of SSIM, achieving the best reconstruction performance among SOTAs. When compared to recent Transformer-based methods such as DAUHST [5], RDLUF [13], DERNN [14], and SSR-L* [52], our CS-Mamba demonstrates notable improvements over these methods, which are 2.06dB, 0.85dB, 0.49dB, and 0.15dB on average, respectively.
We also conduct comparisons of our CS-Mamba-S with previous Transformer-based methods DAUHST [5] and DERNN [14]. As shown in Table 2, we not only achieve better performance in PSNR and SSIM compared to DAUHST and DERNN but also obtain a more lightweight model. Specifically, the CS-Mamba has 0.08 (M) fewer parameters than the DERNN and 13.73 (G) fewer in FLOPs, which means the smallest parameter count and Computational cost among recent deep unfolding methods. Furthermore, we evaluated the denoiser network (Cross-Scan SSM Block) as an independent end-to-end model, comparing it against previous end-to-end Transformer methods, as presented in Table 1. Specifically, the Spatial-Spectral SSM is 2.4 times faster than MST-S and 4.5 times faster than CST-S, while having 0.12 (M) fewer parameters than the MST-S, which means the smallest parameter count.
Except for quantitative comparisons, we also visualize the results of simulation experiments. We choose one of 28 channels from Scene 2, against the simulation results obtained from eleven SOTA approaches. Visually, as illustrated in Fig. 6, our method yields smoother and cleaner details. The experimental results indicate the capability of our method to generate high-quality HSIs.
4.3 Real Data Results
We apply our model and the proposed masked strategy to reconstruct real HSIs, retraining on the CAVE [38] and KAIST [10] datasets and simulating real conditions with 11-bit shot noise. Fig. 7 compares reconstructions from two real scenes using our original CS-Mamba model, the masked version, and four state-of-the-art methods. When enlarging the image results, we can clearly observe that the original CS-Mamba output exhibits a peculiar grid-like noise around the shoulders and face of the subject. In contrast, our mask method effectively mitigates this issue, producing a more visually pleasing effect compared to previous approaches, achieving smoother and more accurate reconstruction.
We further conducted a separate comparison between the original model and the model employing the masked strategy in Scene 1. As shown in Fig. 4, the original model exhibits some wrinkles and artifacts on the raised areas of the petals, with edges becoming blurred, and the stem displaying certain noise patterns. In contrast, the ground truth image presents smooth elevations. This issue is significantly alleviated with our masked method, obtaining a marked improvement in visual quality. Additional experiments will be included in the supplementary material to further substantiate the validity of our approach and conclusions, while Visual comparisons here sufficiently demonstrate the effectiveness of our masked training method.
Refer to caption Real Measurement Refer to caption Ground Truth Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption GAP-Net DGSMP DiffSCI DAUHST Ours (3stg) Ours (Mask)
Figure 7: Visual comparison of SCI reconstruction methods on real $Scene$ 3 at wavelength 648.0nm.
4.4 Ablation Study
Table 4: Ablation study of CS-Mamba (3stg)’s components
Architecture Params (M) PSNR SSIM
Baseline w/o SSM 0.37 37.41 0.964
with LE-SSM 0.77 38.83 0.975
with CS-SSM 0.57 37.88 0.968
with LE + CS SSM 0.97 39.10 0.977
Break-down Ablation. We conduct an ablation study by removing the LE-SSM block and CS-SSM block from the CS-Mamba network. The experiment results are shown in Table 4. Our goal is to explore the contributions of each mechanism to the results. The baseline barely adopts the parameter estimation, U-shaped framework, and GDFFN block without any SSM modules. Since [14] has demonstrated the effectiveness of GDFFN block, we don’t conduct an ablation study for it.
When we respectively employ LE-SSM and CS-SSM blocks, our model achieves 1.42 dB and 0.47 dB improvements in PSNR over the baseline, respectively. Employing both mechanisms simultaneously can further enhance the model’s performance. Here we need to point out that, in consideration of maintaining lightweight, we only utilized a single-direction scan in CS-SSM. In fact, we could also deploy a four-direction scanning similar to LE-SSM. However, each scan direction implies additional computational costs so we decided not to adopt this way.
Table 5: Comparisons of CS mechanism and common scanning.
Architecture PSNR SSIM
w/o Spectral SSM 38.83 0.975
with CS Mechanism 39.10 0.977
with Common Scanning 38.95 0.975
Cross-Scan Mechanism. To highlight the advantages of our cross-scan mechanism over common spectral scanning, we compared model performance under both approaches in our CS-Mamba (3stg) network. As shown in Table 5, common scanning provided a 0.12 dB improvement, whereas the proposed across-scanning mechanism yielded a significant 0.27 dB increase. Both methods were tested under single-scan conditions, with our method excelling due to its enhanced focus on adjacent bands and pixels, effectively capturing local features across spatial and spectral dimensions. These results underscore the effectiveness of our across-scanning approach.
Table 6: Comparisons of LE mechanism and global scanning.
Architecture PSNR SSIM
w/o Spatial SSM 37.88 0.968
with LE Machenism 38.83 0.975
with Global Scanning 38.66 0.974
Local-Enhancement Mechanism. We also conducted an ablation study of our LE-SSM mechanism in our CS-Mamba (3stg) to assess its impact. As shown in Table 6, starting from a baseline model with a PSNR of 37.88 dB (without spatial SSM), incorporating global four-direction scanning improved PSNR by 0.95 dB and SSIM by 0.007, demonstrating effective global spatial context modeling. Applying our LE mechanism along the spatial dimension further improved PSNR by 0.17 dB compared to global scanning, without introducing additional parameters or computational costs, unlike methods using CNN layers or spatial attention [24]. This enhancement highlights the significance of capturing local features in HSIs and validates the effectiveness of our LE mechanism.
5 Conclusion
In this paper, we propose a novel cross-scan SSM mechanism that more fully leverages the spectral and spatial local features of HSI images, achieving improved performance compared to the common SSM mechanism. Additionally, we identify and elaborate on noise issues encountered in real-world scene reconstruction, presenting a masked training method designed to address these challenges. This approach significantly mitigates the degradation problems in real-scene reconstruction that previous methods failed to resolve, enhancing the robustness of the model. Experiments in the main text and supplementary material validate the effectiveness of our method.